OpenStudy (anonymous):

For every integer \( n\ge 0\), show that \[ n^5 -n \] is divisible by 5. This problem is a tutorial about induction. Notice that the statement is true for n=0. Suppose it is true for n. Let us show it for n+1. Notice that \[ (n+1)^5-(n+1)=n^5+5 n^4+10 n^3+10 n^2+4 n=n^5+5 n^4+10 n^3+\\ 10 n^2+5 n-n= \\ (n^5-n)+5 n^4+10 n^3+10 n^2+5 n \] It is clear that the second side is divisible by 5.

OpenStudy (anonymous):

Missing a vital part of a proof by induction: a conclusion.

OpenStudy (anonymous):

sometimes I just find induction so magical

ganeshie8 (ganeshie8):

nice :) division algorithm also gives a neat proof : n^5-n = n(n^4-1) = n(n^2-1)(n^2+1) = (n-1)n(n+1)(n^2+1) clearly it is divisible by 5 for n = 5k, 5k+1, 5k+4 for\(n = 5k \pm 2\), 5 | (n2+1) ending the proof

OpenStudy (anonymous):

That is a nice proof @ganeshie8